A certain dye absorbs 4530 A and fluoresces at 5080 A these being wavelengths of max. absorption that under conditions 47% of the absorbed energy is emitted. Calculate ratio of quanta emitted to the number absorbed.

Grade:11

5 Answers

Pratham Ashish
14 years ago

let no. of femitted quanta = n1

& of absorbed = n2

frequency of emitted quanta  f1 =  c/ 5080* 10^-10

energy e1  =  h*f1 = h*  c/ 5080* 10^-10

frequency of absorbed quanta f2 = c/ 4530* 10^-10

energy e2 =   h*  c/ 4530* 10^-10

   total energy of emitted quanta =  0.47 0f  absorbed quanta

hence,

 n1* h*  c/ 5080* 10^-10 =   0.47*n2* h*  c/ 4530* 10^-10

n1/5080 = 0.47 * n2 /4530

n1/n2 =  0.527

Ramesh V
14 years ago

Quantum efficiency(Q.E) is defined as ratio of quanta emitted to the number absorbed

it is expressed in ratio of energies

quanta absorbed on dye : E = hc/lamda = 1242 eV-nm /453 nm = 2.742 eV

quanta exhibited/undergo fluorescence on dye : E = hc/lamda = 1242 eV-nm /508 nm = 2.445 eV

 

Q.E = 0.47*2.742 / 2.445

       = 0.527

or 52.7 % efficient

I'm not sure if the method was correct, plz verify with someone else

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and
we will get you the answer and detailed solution very quickly.We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch

Shounak yadav
5 years ago
E(absorbed)=1242/453=2.742 
E(energy of photon)=1242/508=2.445
47% of absorbed energy is re-emitted out:
47/100*2.742
no. of photons emitted=energy beingemitted/energy of 1 photon
=47/100*2.742/2.444=0.527
Yash Chourasiya
3 years ago
Hello Student

Energy of light absorbed in an photon= hc/λabsorbed​
Letn1​photons are absorbed
Total energy absorbed= n1​hc/λabsorbed
Now E of light re-emitted out in one photon= hc/λemitted
Letn2​photons are re-emitted then
Total energy re-emitted out= n2​hc/λemitted
Eabsorbed​×47/100 = Ere−emittedout​
hc/λabsorbed​× n1 ​× 47/100 = n2 × hc/λemitted​
n2​/n1​ = 47/100 × λemitted​ /λabsorbed​
= (47/100) × (5080/4530)
n2​/n1​ = 0.527

I hope this solution will help you.
ankit singh
3 years ago
A certain dye absorbs 4530A∘ and fluoresence at 5080A∘ these being wavelength of maximum absorption that under given condition 47% of the absorbed energy is emitted. Calculate the ratio of the no of quanta emitted to the number absorbed.